OpenStudy (anonymous):

How would you simplify the complex fraction? n-7/n^2+10n+24/n-1/n+6 I don't know how to do fractions..Thank you

OpenStudy (campbell_st):

well dividing by a fraction requires you to flip the 2nd fraaction and multiply \[\frac{n - 7}{n^2 + 10n + 24} \times \frac{n + 6}{n - 1}\] factor the quadratic, remove common factors then multiply

OpenStudy (jango_in_dtown):

Is it \[n^2+10n+24=(n+4)(n+6)\]

OpenStudy (jango_in_dtown):

now you can simplify i hope

OpenStudy (anonymous):

Thank you!!!